# Minimality of free-boundary axial hyperplanes in high dimensional circular cones via calibration

### v1

@@ Abstract.
Consider an $(n+1)$-dimensional circular cone. Using a calibration argument, we prove that if $n\geq 4$ and the aperture of the cone is sufficiently large, the intersection of the cone with an axial hyperplane is area-minimizing with respect to free-boundary variations inside the cone.
@@
relative perimeter, calibration, free boundary, minimal cones
@@
The author thanks Gian Paolo Leonardi for constructive discussions on the topics of the paper. The author has been supported by GNAMPA (INdAM) Project 2023: “Esistenza e proprietà fini di forme ottime”; GNAMPA (INdAM) Project 2025: “Structures of sub-Riemannian hypersurfaces in Heisenberg groups”.
@@ 1. Introduction
From a mathematical point of view, given an open set $\Omega\subset\mathbb{R}^{n+1}$, a well-established internal regularity theory for minimizers of (1.1) is available in the literature. In fact, let $E$ be a minimizer of (1.1) and indicate by $\partial^{\ast}E$ its reduced boundary. Letting $M:=\overline{\partial^{\ast}E\cap\Omega}$, one can show that $M\cap\Omega$, that is, the internal portion of $M$, is formed by the union of a smooth hypersurface and a singular set $\Sigma$ whose Hausdorff dimension $\mathop{\rm dim}_{{\mathcal{H}}}\Sigma$ does not exceed the critical value n7n-7, being in particular nonempty only when $n\geq 7$; see [6] and the references therein.
Much less is known regarding the boundary regularity for minimizers of (1.1), that is, the regularity of the free-boundary interface $M\cap\partial\Omega$. One of the first general regularity results, which holds for smooth containers $\Omega\subset\mathbb{R}^{3}$, is due to Taylor [35], who proved that $M\cap\partial\Omega$ does not exhibit singularities. This result has been generalized much later by De Philippis and Maggi in [6] for minimizers of anisotropic capillarity functionals. In the isotropic case, which includes the energy functional (1.1), their result ensures that, if $\partial\Omega$ is $C^{1,1}$, then $M$ is the union of a $C^{1,1/2}$ hypersurface with boundary and a singular set $\Sigma$ with the property that $\mathop{\rm dim}_{{\mathcal{H}}}(\Sigma\cap\partial\Omega)\leq n-3$. Further improvements on the Hausdorff dimension of $\Sigma\cap\partial\Omega$ have been provided in [4] when $\gamma$ is close to either 0, 11 or 1-1. See also [32] for stability and rigidity properties of singular capillary cones arising in the capillarity setting.
@@
The common denominator of the boundary regularity results mentioned above, both for $\gamma\neq 0$ and $\gamma=0$, is that the container $\Omega$ must be smooth. The literature concerning boundary regularity becomes sparser when one looks for results involving nonsmooth containers $\Omega$. In this situation, the main available studies typically consider specific nonsmooth domains or classes of domains. This is the case, for example, in the series of works [16, 17, 18, 19] by Hildebrandt and Sauvigny, where the authors perform a thorough study of minimal surfaces in wedges (i.e., three-dimensional domains obtained as the intersection of two half-spaces); in particular, in [19, p. 73], they prove a regularity result for free-boundary minimal graphs on a plane orthogonal to the wedge.
The latter result has been recently generalized by Edelen and Li in [8]. In that work, they prove a boundary regularity theorem in the same spirit as Grüter-Jost [15], holding for a varifold $M$ under the assumption that $\Omega$ is locally polyhedral and that $M$ is close to a horizontal plane (for instance, if $\Omega=W^{2}\times\mathbb{R}^{n-1}$ is a wedge, horizontal planes are those containing $W^{2}\times\{0\}$). In the same work, Edelen and Li also prove that $\mathop{\rm dim}_{{\mathcal{H}}}(\Sigma\cap\partial\Omega)\leq n-2$, with a sharper estimate holding when the dihedral angles of their cone models are nonobtuse; see [8, Theorem 1.2]. In the setting of their polyhedral domains, the singular set $\Sigma$ is defined as the complement of the points around which the interface $M$ is the graph of a $C^{1,\beta}$ function over a horizontal plane.
In this paper, we prove that in dimensions $n\geq 4$ and for circular cones with sufficiently large aperture, axial hyperplanes are area-minimizing with respect to free-boundary variations. This result fits naturally into some very recent contributions by the author in collaboration with Leonardi [25, 23]. In the latter, we establish a boundary monotonicity formula that holds for almost minimizers of the relative perimeter under minimal regularity assumptions on $\partial\Omega$, and we then employ this formula to prove a minimizing-cone property for a blow-up sequence at a boundary point of $\Omega$. In [25], the authors prove a Vertex-skipping theorem, which states that the boundary of any almost minimizer of the relative perimeter $E$ in a convex domain $\Omega\subset\mathbb{R}^{3}$ cannot contain isolated singularities (say, vertices) of $\partial\Omega$. We highlight that, after a dimension-reduction argument, this result combined with [8, Theorem 1.1] implies the bound $\mathop{\rm dim}_{{\mathcal{H}}}(\Sigma\cap\partial\Omega)\leq n-3$, improving the estimate given in [8, Theorem 1.2 (1)]. The study of the behavior of $E$ near a singularity of the container is also relevant from the perspective of free-boundary minimal surfaces, since the interface $M=\overline{\partial^{\ast}E\cap\Omega}$ is a free-boundary minimal hypersurface in $\Omega$. We refer to [3] and the references therein for background on free-boundary minimal surfaces.
@@
Note that the opening angle $\alpha_{\lambda}\in(0,\pi)$ of $\Omega_{\lambda}$ is related with $\lambda$ by $\lambda=\cot(\alpha_{\lambda}/2)$. Then, if $n=2$, a hyperplane of $\mathbb{R}^{3}$ containing the axis of the cone (up to rotations, $x_{1}=0$) is unstable in $\Omega_{\lambda}$ for every $\lambda>0$. This is in accordance with the aforementioned Vertex-skipping theorem. In partial contrast, when $n\geq 3$, we show that there exists a threshold $\lambda^{\ast}(n)>0$ such that if $0<\lambda\leq\lambda^{\ast}(n)$, i.e., the circular cone has a sufficiently large opening angle $\alpha\geq\alpha_{\lambda^{\star}(n)}$, the axial hyperplane is stable in $\Omega_{\lambda}$, whereas for $\lambda>\lambda^{\ast}(n)$ it is unstable. In this paper, we intend to analyze whether, for $n\geq 3$ and $\lambda\leq\lambda^{\ast}(n)$, an axial hyperplane is merely stable or even area-minimizing in $\Omega_{\lambda}$. We will provide a basically positive answer for $n\geq 4$.
We say that a measurable set $E\subset\mathbb{R}^{n+1}$ has locally finite perimeter provided $P(E;U)<+\infty$ for every $U\subset\!\subset\mathbb{R}^{n+1}$, see [12, 26]. For $(x,t)\in\mathbb{R}^{n+1}$ and $r>0$, we denote by $B_{r}(x,t)$ the ball of radius $r$ centered at $(x,t)$ (with $B_{r}:=B_{r}(0,0)$). Our main result is the following:
@@
$P(E;\Omega_{\lambda}\cap B_{R})\leq P(F;\Omega_{\lambda}\cap B_{R}).$
@@
We emphasize that the above theorem provides a definitive counterexample to the validity of a Vertex-skipping theorem such as [25, Theorem 1.1] in higher dimensions, but only when $n\geq 4$. The case $n=3$ is a limiting situation where, in view of [24], the portion of hyperplane $H_{\lambda}:=\partial E\cap\Omega_{\lambda}$ is stable with respect to free-boundary variations in $\Omega_{\lambda}$, $\lambda\geq\lambda^{\ast}$, but our minimality result, Theorem 1.1, does not apply. We also remark that, despite $\Omega_{\lambda}$ not being a polyhedral cone, Theorem 1.1 appears to suggest that the dimensional bound for the singular set of the free-boundary proved in [8, Theorem 1.2 (1)] cannot be improved beyond the threshold n4n-4. In fact, when $n=4$ and $\lambda\leq 1/2\sqrt{2}$, the result above implies that $\partial E\cap\partial\Omega_{\lambda}=\{0\}$, and so $0\in\Sigma\cap\partial\Omega_{\lambda}$.
@@
The proof of Theorem 1.1 is based on a calibration argument. The calibration method was originally introduced by Bombieri, De Giorgi, and Giusti in their celebrated work [2] to prove the minimality of Simon’s cone in $\mathbb{R}^{8}$. Since then, several authors have contributed by proposing new versions of the original method [21, 10, 29, 5, 7] or by developing various generalizations to different contexts, such as the Steiner tree problem [27, 28, 33] and the Mumford-Shah functional [1, 30].
@@
Here, the calibration method is adapted to the free-boundary setting in the sense that we require the calibrating vector field $Z$ to be tangent to $\partial\Omega_{\lambda}$. This turns out to be crucial for carrying out the calibration argument, as it allows us to exclude the contribution coming from the area of $(E\Delta F)\cap\partial\Omega_{\lambda}$. We point out that the vector field $Z$ will not be defined in the entire closed cone $\overline{\Omega}_{\lambda}$, but only outside a $2$-dimensional subset of $\overline{\Omega}_{\lambda}$. This will force us to perform an approximation argument to apply the Divergence Theorem on $(E\Delta F)\cap\Omega_{\lambda}$, see Section 5.
@@
The existence of the vector field $Z$ will be shown in Section 4, and its construction will be carried out in two steps. Let $S_{\lambda}:=\partial\Omega_{\lambda}\setminus\{0\}$, and denote by $e_{1},\dots,e_{n},e_{t}$ the canonical basis of $\mathbb{R}^{n+1}$. We begin by calibrating the (n1)(n-1)-dimensional surface $S^{0}_{\lambda}:=H_{\lambda}\cap S_{\lambda}$ in $S_{\lambda}$; that is, we build a divergence-free vector field $Y$, defined on a suitable subset $S_{\lambda}^{\prime}$ of $S_{\lambda}$, such that $Y|_{S^{0}_{\lambda}}=e_{1}$ and $|Y|\leq 1$. In a slightly different formulation via differential forms, this was essentially done by Morgan in [31], using ideas from Lawlor’s PhD thesis [20]. Second, we extend the calibration obtained in the previous step to (a suitable portion of) $\Omega_{\lambda}$ by vertical projection onto $S_{\lambda}$: we will show that, due to the structure of the metric on $S_{\lambda}$, the vector field $Z$ obtained in this way is, in particular, divergence-free in its domain.
@@
In Section 2 we provide some preliminary notation and facts concerning differentiable manifolds (especially Riemannian ones) and sets of finite perimeter. In Section 3 we introduce further notation and perform some basic computations needed in the subsequent parts of the paper. Then, Section 4 is devoted to the construction of the calibration $Z$, whose main properties are collected in Theorem 4.1. Finally, in Section 5 we present the calibration argument; namely, we show how the existence of the vector field $Z$ from Theorem 4.1 can be used to demonstrate Theorem 1.1.
@@ 2. Preliminary notions and basic notation
It can be shown that $\star\circ\star=\star^{2}$ coincides with the multiplication by $(-1)^{k(n-k)}$, and that $\star\nu=1$. In particular, when k=n1k=n-1, the operator $\star$ allows to associate to every (n1)(n-1)-form $\omega\in\mathcal{A}^{n-1}(M)$ a $1$-form $\star\omega\in\mathcal{A}^{1}(M)$. Given a vector field $X\in\mathfrak{X}(M)$, denote by $i_{X}\nu\in\mathcal{A}^{n-1}(M)$ the contraction between $X$ and the volume form $\nu$. We have the following technical Lemma (see [22, p. 52]).
@@ 2.4. Locally finite perimeter sets under set operations
If $E$ is a locally finite perimeter set, we indicate by $\partial^{\ast}E$ its reduced boundary and by $\nu_{E}$ the inner unit normal to its reduced boundary [26]. Assume now that $E$ and $F$ have locally finite perimeter. We denote by $Q(E,F)^{\pm}$ the set
$Q(E,F)^{\pm}:=\{(x,t)\in\partial^{\ast}E\cap\partial^{\ast}F\,:\,\nu_{E}(x,t)=\pm\nu_{F}(x,t)\}.$
It is well known that also $E\cap F$ and $E\setminus F$ have locally finite perimeter and, up to ${\mathcal{H}}^{n}$-negligible sets, one has (see [26, Theorem 16.3])

### v2

@@ Abstract.
Consider an $(n+1)$-dimensional circular cone with opening angle $\alpha\in(0,\pi)$. Using a free-boundary adaptation of the classical calibration method, we prove that, for $n\geq 4$, there exists a threshold $\bar{\alpha}(n)\in(0,\pi)$ such that if $\alpha\geq\bar{\alpha}(n)$, that is, the cone is wide enough, the intersection of the cone with an axial hyperplane is area-minimizing with respect to free-boundary variations inside the cone. This provides a counterexample to a recent Vertex-skipping Theorem proved by the author in collaboration with G.P. Leonardi, at least for $n\geq 4$.
@@
relative perimeter, calibration, free-boundary, minimal cones
@@
The author thanks Gian Paolo Leonardi for constructive discussions on the topics of the paper. The author has been supported by GNAMPA (INdAM) Project 2025: “Structures of sub-Riemannian hypersurfaces in Heisenberg groups”; GNAMPA (INdAM) Project 2026: “Variational, Geometric, and Analytic Perspectives on Regularity”.
@@ 1. Introduction
From a mathematical point of view, given an open set $\Omega\subset\mathbb{R}^{n+1}$, a well-established internal regularity theory for minimizers of (1.1) is available in the literature. In fact, let $E$ be a minimizer of (1.1) and indicate by $\partial^{\ast}E$ its reduced boundary. Letting $M:=\overline{\partial^{\ast}E\cap\Omega}$, it is known that $M\cap\Omega$, that is, the internal portion of $M$, is formed by the union of a smooth hypersurface and a singular set $\Sigma$ whose Hausdorff dimension $\mathop{\rm dim}_{{\mathcal{H}}}\Sigma$ does not exceed the critical value n7n-7, being in particular nonempty only when $n\geq 7$; see [7] and the references therein.
Much less is known regarding the boundary regularity for minimizers of (1.1), that is, the regularity of the free-boundary interface $M\cap\partial\Omega$. One of the first general regularity results, which holds for smooth containers $\Omega\subset\mathbb{R}^{3}$, is due to Taylor [41], who proved that $M\cap\partial\Omega$ does not exhibit singularities. This result has been generalized much later by De Philippis and Maggi in [7] for minimizers of anisotropic capillarity functionals. In the isotropic case, which includes the energy functional (1.1), their result ensures that, if $\partial\Omega$ is $C^{1,1}$, then $M$ is the union of a $C^{1,1/2}$ hypersurface with boundary and a singular set $\Sigma$ with the property that $\mathop{\rm dim}_{{\mathcal{H}}}(\Sigma\cap\partial\Omega)\leq n-3$. Further improvements on the Hausdorff dimension of $\Sigma\cap\partial\Omega$ have been provided in [5] when $\gamma$ is close to either 0, 11 or 1-1. See also [9] for a regularity result on $\Lambda$-minimizers of (1.1) under the assumption that $\Omega$ is a half-space and the wet region of competitors is constrained within an open subset of $\partial\Omega$; and [38] for stability and rigidity properties of singular capillary cones.
@@
The common denominator of the boundary regularity results mentioned above, both for $\gamma\neq 0$ and $\gamma=0$, is that the container $\Omega$ must be smooth. The literature concerning boundary regularity becomes sparser when one looks for results involving nonsmooth containers $\Omega$. In this situation, the main available studies typically consider specific nonsmooth domains or classes of domains. This is the case, for example, in the series of works [18, 19, 20, 21] by Hildebrandt and Sauvigny, where the authors perform a thorough study of minimal surfaces in wedges (i.e., three-dimensional domains obtained as the intersection of two half-spaces); in particular, in [21, p. 73], they prove a regularity result for free-boundary minimal graphs on a plane orthogonal to the wedge.
The latter result has been recently generalized by Edelen and Li in [10]. In that work, they prove a boundary regularity theorem in the same spirit as Grüter-Jost [17], holding for a varifold $M$ under the assumption that $\Omega$ is locally polyhedral and that $M$ is close to a horizontal plane (for instance, if $\Omega=W^{2}\times\mathbb{R}$ is a wedge, horizontal planes are those containing $W^{2}\times\{0\}$). In the same work, Edelen and Li also prove that $\mathop{\rm dim}_{{\mathcal{H}}}(\Sigma\cap\partial\Omega)\leq n-2$, with a sharper estimate holding when the dihedral angles of their cone models are nonobtuse; see [10, Theorem 1.2]. In the setting of their polyhedral domains, the singular set $\Sigma$ is defined as the complementary set of the points around which the interface $M$ is the graph of a $C^{1,\beta}$ function over a horizontal plane.
In this paper, we prove that for $n\geq 4$ and for circular cones with sufficiently large aperture, axial hyperplanes are area-minimizing with respect to free-boundary variations. This result fits naturally into some very recent contributions by the author in collaboration with Leonardi [28, 29] concerning boundary regularity in nonsmooth domains. In the latter, we establish a boundary monotonicity formula that holds for almost minimizers of the relative perimeter under mild regularity assumptions on $\partial\Omega$, which is then employed to prove a minimizing-cone property for a blow-up sequence at a boundary point of $\Omega$. In [28], the authors prove a Vertex-skipping Theorem, which states that the boundary of any almost minimizer of the relative perimeter $E$ in a convex domain $\Omega\subset\mathbb{R}^{3}$ cannot contain isolated singularities (say, vertices) of $\partial\Omega$. We highlight that, after a dimension-reduction argument, this result combined with [10, Theorem 1.1] implies the bound $\mathop{\rm dim}_{{\mathcal{H}}}(\Sigma\cap\partial\Omega)\leq n-3$, improving the estimate given in [10, Theorem 1.2 (1)].
@@
Note that the opening angle $\alpha_{\lambda}\in(0,\pi)$ of $\Omega_{\lambda}$ is related with $\lambda$ by $\lambda=\cot(\alpha_{\lambda}/2)$. Then, if $n=2$, a hyperplane of $\mathbb{R}^{3}$ containing the axis of the cone (up to rotations, $x_{1}=0$) is unstable in $\Omega_{\lambda}$, for every $\lambda>0$. This is in accordance with the aforementioned Vertex-skipping Theorem. In partial contrast, when $n\geq 3$, we show that there exists a threshold $\lambda^{\ast}(n)>0$ such that if $0<\lambda\leq\lambda^{\ast}(n)$, i.e., the circular cone has a sufficiently large opening angle $\alpha\geq\alpha^{\ast}(n):=\alpha_{\lambda^{\star}(n)}$, the axial hyperplane is stable in $\Omega_{\lambda}$, whereas for $\lambda>\lambda^{\ast}(n)$, or equivalently if $\alpha<\alpha^{\ast}(n)$, it is unstable. In this paper, we intend to analyze whether, for $n\geq 3$ and $0<\lambda\leq\lambda^{\ast}(n)$, an axial hyperplane is merely stable or even area-minimizing in $\Omega_{\lambda}$. We will provide a basically positive answer for $n\geq 4$.
We say that a measurable set $E\subset\mathbb{R}^{n+1}$ has locally finite perimeter provided $P(E;U)<+\infty$ for every $U\subset\!\subset\mathbb{R}^{n+1}$, see [14, 31]. For $(x,t)\in\mathbb{R}^{n+1}$ and $r>0$, we denote by $B_{r}(x,t)$ the ball of radius $r$ centered at $(x,t)$ (setting $B_{r}:=B_{r}(0,0)$). Our main result is the following:
@@
$P(E;\Omega_{\lambda}\cap B_{R})\leq P(F;\Omega_{\lambda}\cap B_{R}).$ (1.3)
In other words, if the opening angle $\alpha$ of the circular cone satisfies $\alpha\geq\bar{\alpha}(n):=\alpha_{\bar{\lambda}(n)}$, then $H_{\lambda}:=\partial E\cap\Omega_{\lambda}$ is a free-boundary area-minimizing surface inside the cone.
@@
We emphasize that the above theorem provides a definitive counterexample to the validity of a Vertex-skipping Theorem such as [28, Theorem 1.1] in higher dimensions, but only when $n\geq 4$. The case $n=3$ is a limiting situation where, in view of [27], provided $0<\lambda\leq\lambda^{\ast}(n)$, $H_{\lambda}$ is stable with respect to free-boundary variations in $\Omega_{\lambda}$, but our minimality result, Theorem 1.1, does not apply. We also remark that, despite $\Omega_{\lambda}$ not being a polyhedral cone, Theorem 1.1 appears to suggest that the dimensional bound for the singular set of the free-boundary proved in [10, Theorem 1.2 (1)] cannot be improved beyond the threshold n4n-4. In fact, when $n=4$ and $0<\lambda\leq 1/2\sqrt{2}$, the result above implies that $\partial E\cap\partial\Omega_{\lambda}=\{0\}$, and so $0\in\Sigma\cap\partial\Omega_{\lambda}$.
@@
The proof of Theorem 1.1 is based on a calibration argument. The calibration method was originally introduced by Bombieri, De Giorgi, and Giusti in their celebrated work [3] to prove the minimality of Simon’s cone in $\mathbb{R}^{8}$. Since then, several authors have contributed by proposing new versions of the original method [25, 12, 34, 30, 22, 4, 23, 6, 8] or by developing various generalizations to different contexts, such as the Steiner tree problem [32, 33, 39] and the Mumford-Shah functional [2, 1, 35]. See [37, §6.5] for a more comprehensive overview on calibrations.
@@
Here, the calibration method is adapted to the free-boundary setting in the sense that we require the calibrating vector field to be tangent to $\partial\Omega_{\lambda}$. Our calibration argument, explained in detail in Section 5, proceeds roughly as follows. For $\lambda>0$, let $S_{\lambda}:=\partial\Omega_{\lambda}\setminus\{0\}$, and denote by $e_{1},\dots,e_{n},e_{t}$ the canonical basis of $\mathbb{R}^{n+1}$. First, we construct a divergence-free vector field $Z$, defined outside a $2$-dimensional subset of $\overline{\Omega}_{\lambda}$, such that $Z|_{H_{\lambda}}=e_{1}$, $|Z|\leq 1$, and $Z|_{S_{\lambda}}$ is tangent to $S_{\lambda}$. Second, we apply the Divergence Theorem to $Z$ on the region $(E\Delta F)\cap\Omega_{\lambda}$. On the one hand, the fact that $Z$ is tangent to $S_{\lambda}$ allows us to disregard the contribution arising from the area of $(E\Delta F)\cap\partial\Omega_{\lambda}$. On the other hand, the condition $\operatorname{div}Z=0$ enables us to compare $P(E;\Omega_{\lambda}\cap B_{R})$ with $P(F;\Omega_{\lambda}\cap B_{R})$. We stress that here, as remarked above, the vector field $Z$ is not defined on the entire closed cone $\overline{\Omega}_{\lambda}$, but only outside a $2$-dimensional subset of $\overline{\Omega}_{\lambda}$. This forces us to perform an approximation argument in order to apply the Divergence Theorem to $(E\Delta F)\cap\Omega_{\lambda}$ and infer the desired inequality (1.3); see Section 5.
@@
The existence of the vector field $Z$ will be exhibited in Section 4, and its construction will be carried out in two steps. We begin by calibrating the (n1)(n-1)-dimensional surface $S^{0}_{\lambda}:=H_{\lambda}\cap S_{\lambda}$ in $S_{\lambda}$; that is, we build a divergence-free vector field $Y$, defined on a suitable subset $S_{\lambda}^{\prime}$ of $S_{\lambda}$ and tangent to $S_{\lambda}$, such that $Y|_{S^{0}_{\lambda}}=e_{1}$ and $|Y|\leq 1$. In a slightly different formulation via differential forms, this was essentially done by Morgan in [36], using ideas from Lawlor’s PhD thesis [24]. Their construction basically reduces the problem to finding a solution to a suitable differential inequality subject to appropriate initial conditions (cf. the proof of Theorem 4.2). Such a solution exists for $n\geq 4$ and $0<\lambda\leq\bar{\lambda}(n)$, thereby explaining the dimensional gap occurring for $n=3$. Second, we extend $Y$ to (a suitable portion of) $\Omega_{\lambda}$ by vertical projection onto $S_{\lambda}$: we will show that, due to the structure of the metric on $S_{\lambda}$, the vector field $Z$ obtained in this way is, in particular, divergence-free in its domain (cf. the proof of Theorem 4.1).
@@
In Section 2 we provide some preliminary notation and notions concerning differentiable manifolds (especially Riemannian ones) and sets of finite perimeter. In Section 3 we introduce further notation and perform some basic computations needed in the subsequent parts of the paper. Then, Section 4 is devoted to the construction of the calibration $Z$, whose main properties are collected in Theorem 4.1. Finally, in Section 5 we present the calibration argument; namely, we show how the existence of the vector field $Z$ from Theorem 4.1 can be used to demonstrate Theorem 1.1.
@@ 2. Preliminary notions and basic notation
It can be shown that $\star\circ\star=\star^{2}$ coincides with the multiplication by $(-1)^{k(n-k)}$, and that $\star\nu=1$. In particular, when k=n1k=n-1, the operator $\star$ allows to associate to every (n1)(n-1)-form $\omega\in\mathcal{A}^{n-1}(M)$ a $1$-form $\star\omega\in\mathcal{A}^{1}(M)$. Given a vector field $X\in\mathfrak{X}(M)$, denote by $i_{X}\nu\in\mathcal{A}^{n-1}(M)$ the contraction between $X$ and the volume form $\nu$. We have the following technical Lemma (see [26, p. 52]).
@@ 2.4. Locally finite perimeter sets under set operations
If $E$ has locally finite perimeter, we indicate by $\partial^{\ast}E$ its reduced boundary and by $\nu_{E}$ the inner unit normal to $\partial^{\ast}E$ [31]. Assume now that $E$ and $F$ have locally finite perimeter. We define the set $Q(E,F)^{\pm}:=\{(x,t)\in\partial^{\ast}E\cap\partial^{\ast}F\,:\,\nu_{E}(x,t)=\pm\nu_{F}(x,t)\}.$ It is well known that $E\cap F$, $E\setminus F$ have also locally finite perimeter and, up to ${\mathcal{H}}^{n}$-negligible sets, one has (see [31, Theorem 16.3])